# Influence Functions for Edge Edits in Non-Convex Graph Neural Networks

### v1

@@ 1 Introduction
In this work, we propose an influence function tailored specifically for GNNs, enabling precise predictions regarding the effects of edge modifications from multiple perspectives. To do so, we overcome the unique challenges of GNNs by deriving the changes in the evaluation function from fundamental principles of calculus. As a result, our function can measure the influence of both edge deletion and insertion, the latter of which has not been explored in previous studies. To extend the influence function to non-convex GNNs, we establish the proximal Bregman response function [3] for node classification. The influence function derived from this response function relies on weaker assumptions that generally hold for non-convex GNNs. Figure 1 shows an example of influence analysis for a barbell graph from two different perspectives. Through the analysis, one can identify that the edges connecting two different clusters have opposite influences on over-squashing and over-smoothing, providing a unifying view on edge importance.
@@ 2.1 Influence function for convex models
When the objective is non-convex, the Hessian can be added with a damping term $\lambda\in\mathbb{R}$, i.e., $\mathbf{H}_{\theta^{*}}+\lambda I$, leading to a positive definite matrix. We provide a complete derivation in Appendix A.
@@ 2.2 Influence function for neural networks
Although the influence function in Equation 4 reliably measures the influence of an input for a linear model with a convex objective, the computation is unreliable in practice with deep neural networks, c.f. [4]. Bae et al. [3] identify the three main sources of unreliability related to the standard practices of neural network training and fine-tuning: 1) In non-convex models, the response function is affected more by parameter initialization than by the influence of the data, as gradient methods approximate the solution. 2) The addition of the damping term $\lambda$ works as an $\ell_{2}$ regularizer of Equation 1 leading to a different response function. 3) The influence function is measured on fully converged parameter $\theta^{*}$, which is not true in practice due to many reasons, such as early stopping and over-fitting mitigation.
@@
For detailed explanations of each term in Equation 5, please refer to Bae et al. [3].
The objective can be further linearized around the model output to simulate the local approximations made in Equation 4, leading to the following influence function [^1] Equation 7 is originally proposed by Teso et al. [29] as an approximation of Equation 4. Bae et al. [3] provide a corresponding response function to the approximation.:
@@
where $h$ and $h_{s}$ denote the model outputs parameterized by $\theta$ and $\theta_{s}$, respectively. $\mathbf{J}_{h\theta_{s}}$ is the Jacobian of the model outputs with respect to the parameters, and $\mathbf{H}_{h_{s}}$ is the Hessian of $\mathcal{L}$ with respect to the outputs. When the loss function is convex with respect to the model outputs, the matrix $\mathbf{J}_{h\theta_{s}}^{\top}\mathbf{H}_{h_{s}}\mathbf{J}_{h\theta_{s}}$ is positive semi-definite. This convexity condition is satisfied by commonly used loss functions, such as cross-entropy and mean squared error. Consequently, for $\lambda>0$, the matrix $\mathbf{J}_{h\theta_{s}}^{\top}\mathbf{H}_{h_{s}}\mathbf{J}_{h\theta_{s}}+\lambda\mathbf{I}$ becomes positive definite, ensuring its invertibility. We provide the complete derivation in Appendix A.
@@ 3 Quantifying the influence of edge edits on GNNs
We focus on analyzing the effect of inserting or deleting a single undirected edge $\{u,v\}$. We first define the reweighted adjacency matrix $A^{\epsilon}_{uv}{}$ such that only the $(u,v)$ and $(v,u)$ entries are updated as $A^{\epsilon}_{uv}{}=A_{uv}+(2\mathbb{I}[\{u,v\}\in\mathcal{E}]-1)N\epsilon$, while all other entries remain unchanged. Setting $\epsilon=-1/N$ corresponds to deleting the edge if it exists, or inserting it otherwise. We denote the edge-reweighted graph as $\mathcal{G}^{\epsilon}=\{\mathcal{V},\mathcal{E},A^{\epsilon}{}\}$, where we omit the target edge $\{u,v\}$ from the adjacency matrix when it is clear from context for notational simplicity. We collectively refer to edge deletions and insertions as edge edits.
@@
To quantify the change in model parameters for non-convex GNNs, we propose a graph-adapted version of the PBRF. The original PBRF, introduced in Equation 5, only quantifies the changes in parameters when the weight of a single data point is modified, inapplicable to our scenario where the weight of an edge changes. To address this challenge, we introduce an edge-edit PBRF that explicitly accounts for changes in node representations caused by an edge edit.
@@
where $\mathbf{G}=\mathbf{J}_{h\theta_{s}}^{\top}\mathbf{H}_{h_{s}}\mathbf{J}_{h\theta_{s}}+\lambda\mathbf{I}$ denotes the generalized Gauss–Newton Hessian with a damping term, and $h$ and $h_{s}$ denote the node representations obtained using parameters $\theta$ and $\theta_{s}$, respectively. A detailed derivation is provided in Appendix B.
@@
To quantify the changes in the evaluation function caused by the modification in the message propagation path, we can further expand the message propagation term in Equation 8. Since $\partial A^{\epsilon}_{ij}/\partial\epsilon=0$ for all $\{i,j\}\neq\{u,v\}$, the message propagation term simplifies as follows:
@@
By substituting Section 3 and Equation 10 into Equation 8 and linearizing around $\epsilon=0$, we obtain a first-order approximation of the influence function that captures both parameter shift and message propagation effects. The resulting influence of an edge edit is given by:
@@
Directly computing $\mathbf{G}^{-1}$ is computationally infeasible for large models. To address this, we approximate the inverse Hessian-vector product $\mathbf{G}^{-1}\nabla_{\theta}f(\theta_{s},\mathcal{G})$ using the LiSSA algorithm [1], a stochastic iterative method. A detailed description is provided in Appendix C.
@@ 4.1 Evaluation functions
Figure 2: Predicted influence vs. actual influence on four-layer GCN. The x-axis represents the predicted influence, the y-axis represents the actual influence, and the dotted line represents the perfect alignment.
@@
We evaluate the prediction performance of our method on three representative graph neural networks: GCN [19], GAT [31], and ChebNet [12]. We compare our approach with GIF, an existing graph influence function [6, 33]. To measure the actual influence, we first train the model on the original graph. We then retrain the GNN on the edge-edited graph by optimizing the minimization objective defined by each response function. Specifically, for our influence function, we fine-tuned the model by minimizing the objective in Equation 9, using the original model parameters as $\theta_{s}$ and setting $\epsilon=-1/N$. For GIF, we retrain the model by minimizing the loss on the edge-edited graph, using the same parameter initialization as the original model. The actual influence is computed as the difference between the evaluation values of the original model and the retrained model.
@@ Results
Figure 2 presents scatter plots for the Cora dataset, where each point shows the actual and predicted influence of a single edge edit. The x-axis represents the predicted influence computed by each method, and the y-axis represents the actual influence. The first row illustrates results from GIF [6, 33], while the second and third rows show results from our proposed method for edge deletions and insertions, respectively. Additional results for other GNNs and datasets are provided in Appendix D.
@@
Table 1 shows the test accuracy when the model is trained and evaluated on edge-edited graphs. Dirichlet energy (DE), over-squashing ($f_{\text{OQ}}$), and validation loss (VL) are used as shorthand in the table. Our edits based on validation loss achieve the best performance across three datasets. This result suggests that editing edges to reduce validation loss can be helpful for improving test accuracy. In contrast, editing edges to improve Dirichlet energy or the over-squashing measurement does not consistently improve test accuracy. This indicates that optimizing these intermediate metrics does not necessarily translate to better predictive performance. Finally, editing edges to reduce validation loss using GIF fails to improve test accuracy, which we attribute to its inaccurate influence estimation on non-convex GNNs.
@@
Figure 3: The relationship between parameter shift effect and message propagation effect defined in Equation 8. The x-axis denotes the parameter shift effect, and the y-axis denotes the message propagation effect.
@@
Figure 3 shows scatter plots comparing these two influences. We observe a low correlation between the message propagation and parameter shift influence across all metrics, while their magnitudes remain comparable. These results underscore the necessity of explicitly measuring the influence of message propagation in estimating the influence of edge edits.
@@ 5 Applications
Analysis of the effect of edge rewiring
Figure 4: Histograms of the estimated influence of edge insertions (blue) and deletions (red) selected by BORF [24], evaluated on a four-layer GCN trained on the Texas dataset. Influences are measured across validation loss, over-squashing, and over-smoothing. Arrows ($\downarrow$/$\uparrow$ / ↑) indicate the desired direction of each measurement (decrease/increase).
Our approach provides multiple analytical perspectives on existing rewiring strategies through various evaluation metrics. Specifically, we analyze BORF [24], which alleviates over-squashing by inserting edges between nodes with negative curvature and reduces over-smoothing by removing edges with positive curvature.
Figure 4 shows the estimated influence of edges selected by BORF from three measurement perspectives. Edge insertions selected by BORF tend to increase the over-squashing measurement, while edge deletions tend to increase the over-smoothing measurement, confirming that BORF effectively targets its intended GNN challenges. However, we also observe unintended side effects: edge insertions tend to worsen the over-smoothing measurement, and neither insertions nor deletions consistently reduce validation loss. Addressing these unintended effects could significantly improve the overall effectiveness and reliability of BORF. We also provide a similar analysis for FoSR [18] in Appendix E.
@@
Table 2 presents the adversarial attack performance of our methods, comparing it to previous influence-function-based methods and other adversarial attack methods, DICE [32] and PRBCD [13], in the white-box scenario. As a result, the GIF shows poor performance, while our method with a validation loss metric outperforms the others. An interesting observation is that, among the three metrics we tested, the validation loss-based attack is more effective than the tailored attack methods.
@@
We demonstrate the influence of edges connecting the node with the same label (homophilic edge) and different labels (heterophilic edge) on GCN. Figure 5 presents a dumbbell plot showing the mean influence of homophilic and heterophilic edges across six datasets, which include three homophilic graphs (Cora, CiteSeer, and PubMed), where edges tend to connect nodes with the same labels, and three heterophilic graphs (Chameleon, Actor, and Squirrel), where edges tend to connect nodes with different labels.
@@ 7 Conclusion
We introduce an enhanced graph influence function that estimates the impact of edge perturbations on model predictions in Graph Neural Networks. Unlike existing graph influence functions, our approach explicitly incorporates message propagation effects and relaxes the convexity assumption, enabling it to be applied to commonly used non-convex GNNs. Additionally, we extend the framework to handle both edge deletions and insertions, broadening its applicability to real-world graph rewiring tasks. Experimental results on various real-world datasets demonstrate that our method provides significantly more accurate estimates of influence than previous methods and enables effective improvements in key target measurements such as validation loss, over-squashing, and over-smoothing. Despite these advances, our method has limitations, particularly in estimating the influence of multiple edge edits (i.e., removing or adding multiple edges and measuring the joint effect) and scalability to deep GNNs. Future research could focus on improving these aspects to enhance the method’s applicability to larger and more complex graph learning tasks.
@@ Appendix A Derivation of the influence function
In this section, we derive the influence function for linear models [20], as shown in Equation 3, and for neural networks [3], as shown in Equation 7. In particular, we reformulate the derivation by Bae et al. [3] to fit our setting.
@@ A.1 Influence function for linear models
We begin by considering the re-weighted objective introduced in Equation 2:
@@
Under these conditions, the Implicit Function Theorem ensures that $\theta_{x^{\prime},y^{\prime},\epsilon}^{*}$ is continuously differentiable with respect to $\epsilon$, and we differentiate Equation 16 using the chain rule:
@@ A.2 Influence function for neural networks
00 0 $=-\frac{1}{N}\sum_{(x,y)\in\mathcal{D}_{\text{train}}}\mathbf{J}_{h^{\theta}_{x}\theta_{s}}^{\top}\nabla_{h^{\theta}_{x}}^{2}\mathcal{L}(h^{\theta_{s}}_{x})\mathbf{J}_{h^{\theta}_{x}\theta_{s}}(\theta^{*}_{\text{lin},x^{\prime},y^{\prime},\epsilon}-\theta_{s})+\lambda(\theta^{*}_{\text{lin},x^{\prime},y^{\prime},\epsilon}-\theta_{s})+\epsilon\,\nabla_{\theta}\mathcal{L}(x^{\prime},\theta_{s}).$ (24)

### v2

@@ 1 Introduction
In this work, we propose an influence function tailored specifically for GNNs, enabling precise predictions regarding the effects of edge modifications from multiple perspectives. To do so, we overcome the unique challenges of GNNs by deriving the changes in the evaluation function from fundamental principles of calculus. As a result, our function can measure the influence of both edge deletion and insertion, the latter of which has not been explored in previous studies. To extend the influence function to non-convex GNNs, we establish the proximal Bregman response function [3] for node classification. The influence function derived from this response function relies on weaker assumptions that generally hold for non-convex GNNs. Figure˜1 shows an example of influence analysis for a barbell graph from two different perspectives. Through the analysis, one can identify that the edges connecting two different clusters have opposite influences on over-squashing and over-smoothing, providing a unifying view on edge importance.
@@ 2.1 Influence function for convex models
When the objective is non-convex, the Hessian can be added with a damping term $\lambda\in\mathbb{R}$, i.e., $\mathbf{H}_{\theta^{*}}+\lambda I$, leading to a positive definite matrix. We provide a complete derivation in Appendix˜A.
@@ 2.2 Influence function for neural networks
Although the influence function in Equation˜4 reliably measures the influence of an input for a linear model with a convex objective, the computation is unreliable in practice with deep neural networks, c.f. [4]. Bae et al. [3] identify the three main sources of unreliability related to the standard practices of neural network training and fine-tuning: 1) In non-convex models, the response function is affected more by parameter initialization than by the influence of the data, as gradient methods approximate the solution. 2) The addition of the damping term $\lambda$ works as an $\ell_{2}$ regularizer of Equation˜1 leading to a different response function. 3) The influence function is measured on fully converged parameter $\theta^{*}$, which is not true in practice due to many reasons, such as early stopping and over-fitting mitigation.
@@
For detailed explanations of each term in Equation˜5, please refer to Bae et al. [3].
The objective can be further linearized around the model output to simulate the local approximations made in Equation˜4, leading to the following influence function [^1] Equation 7 is originally proposed by Teso et al. [31] as an approximation of Equation 4. Bae et al. [3] provide a corresponding response function to the approximation.:
@@
where $h$ and $h_{s}$ denote the model outputs parameterized by $\theta$ and $\theta_{s}$, respectively. $\mathbf{J}_{h\theta_{s}}$ is the Jacobian of the model outputs with respect to the parameters, and $\mathbf{H}_{h_{s}}$ is the Hessian of $\mathcal{L}$ with respect to the outputs. When the loss function is convex with respect to the model outputs, the matrix $\mathbf{J}_{h\theta_{s}}^{\top}\mathbf{H}_{h_{s}}\mathbf{J}_{h\theta_{s}}$ is positive semi-definite. This convexity condition is satisfied by commonly used loss functions, such as cross-entropy and mean squared error. Consequently, for $\lambda>0$, the matrix $\mathbf{J}_{h\theta_{s}}^{\top}\mathbf{H}_{h_{s}}\mathbf{J}_{h\theta_{s}}+\lambda\mathbf{I}$ becomes positive definite, ensuring its invertibility. We provide the complete derivation in Appendix˜A.
@@ 3 Quantifying the influence of edge edits on GNNs
We focus on analyzing the effect of inserting or deleting a single undirected edge $\{u,v\}$. The derivation naturally generalizes to multiple edge edits, which are provided in Appendix˜C. We first define the reweighted adjacency matrix $A^{\epsilon}_{uv}{}$ such that only the $(u,v)$ and $(v,u)$ entries are updated as $A^{\epsilon}_{uv}{}=A_{uv}+(2\mathbb{I}[\{u,v\}\in\mathcal{E}]-1)N\epsilon$, while all other entries remain unchanged. Setting $\epsilon=-1/N$ corresponds to deleting the edge if it exists, or inserting it otherwise. We denote the edge-reweighted graph as $\mathcal{G}^{\epsilon}=\{\mathcal{V},\mathcal{E},A^{\epsilon}{}\}$, where we omit the target edge $\{u,v\}$ from the adjacency matrix when it is clear from context for notational simplicity. We collectively refer to edge deletions and insertions as edge edits.
@@
To quantify the change in model parameters for non-convex GNNs, we propose a graph-adapted version of the PBRF. The original PBRF, introduced in Equation˜5, only quantifies the changes in parameters when the weight of a single data point is modified, inapplicable to our scenario where the weight of an edge changes. To address this challenge, we introduce an edge-edit PBRF that explicitly accounts for changes in node representations caused by an edge edit.
@@
where $\mathbf{G}=\mathbf{J}_{h\theta_{s}}^{\top}\mathbf{H}_{h_{s}}\mathbf{J}_{h\theta_{s}}+\lambda\mathbf{I}$ denotes the generalized Gauss–Newton Hessian with a damping term, and $h$ and $h_{s}$ denote the node representations obtained using parameters $\theta$ and $\theta_{s}$, respectively. A detailed derivation is provided in Appendix˜B.
@@
To quantify the changes in the evaluation function caused by the modification in the message propagation path, we can further expand the message propagation term in Equation˜8. Since $\partial A^{\epsilon}_{ij}/\partial\epsilon=0$ for all $\{i,j\}\neq\{u,v\}$, the message propagation term simplifies as follows:
@@
By substituting Section˜3 and Equation˜10 into Equation˜8 and linearizing around $\epsilon=0$, we obtain a first-order approximation of the influence function that captures both parameter shift and message propagation effects. The resulting influence of an edge edit is given by:
@@
Directly computing $\mathbf{G}^{-1}$ is computationally infeasible for large models. To address this, we approximate the inverse Hessian-vector product $\mathbf{G}^{-1}\nabla_{\theta}f(\theta_{s},\mathcal{G})$ using the LiSSA algorithm [1], a stochastic iterative method. A detailed description is provided in Appendix˜D.
@@ 4.1 Evaluation functions
Figure 2: Predicted influence versus actual influence on a four-layer GCN. The x-axis represents the predicted influence, the y-axis represents the actual influence, and the dotted line represents the perfect alignment.
@@
We evaluate the prediction performance of our method on three representative graph neural networks: GCN [20], GAT [33], and ChebNet [13]. We compare our approach with GIF, an existing graph influence function [7, 35]. To measure the actual influence, we first train the model on the original graph. We then retrain the GNN on the edge-edited graph by optimizing the minimization objective defined by each response function. Specifically, for our influence function, we fine-tuned the model by minimizing the objective in Equation˜9, using the original model parameters as $\theta_{s}$ and setting $\epsilon=-1/N$. For GIF, we retrain the model by minimizing the loss on the edge-edited graph, using the same parameter initialization as the original model. The actual influence is computed as the difference between the evaluation values of the original model and the retrained model.
@@ Results
Figure˜2 presents scatter plots for the Cora dataset, where each point shows the actual and predicted influence of a single edge edit. The x-axis represents the predicted influence computed by each method, and the y-axis represents the actual influence. The first row illustrates results from GIF [7, 35], while the second row shows results from our proposed method for edge deletions and insertions, respectively. Additional results for other GNNs and datasets are provided in Appendix˜E.
@@
Table˜1 shows the test accuracy when the model is trained and evaluated on edge-edited graphs. Dirichlet energy (DE), over-squashing ($f_{\text{OQ}}$), and validation loss (VL) are used as shorthand in the table. Our edits based on validation loss achieve the best performance across three datasets. This result suggests that editing edges to reduce validation loss can be helpful for improving test accuracy. In contrast, editing edges to improve Dirichlet energy or the over-squashing measurement does not consistently improve test accuracy. This indicates that optimizing these intermediate metrics does not necessarily translate to better predictive performance. Finally, editing edges to reduce validation loss using GIF fails to improve test accuracy, which we attribute to its inaccurate influence estimation on non-convex GNNs.
@@
Figure 3: The relationship between parameter shift effect and message propagation effect defined in Equation˜8. The x-axis denotes the parameter shift effect, and the y-axis denotes the message propagation effect.
@@
Figure˜3 shows scatter plots comparing these two influences on the Cora dataset. We observe a low correlation between the message propagation and parameter shift influence across all metrics, while their magnitudes remain comparable. These results underscore the necessity of explicitly measuring the influence of message propagation in estimating the influence of edge edits.
Influence estimation for multiple edge edits
Figure 4: Predicted influence versus actual influence on a four-layer GCN under varying numbers of inserted edges.
We evaluate the accuracy of the proposed influence function in predicting the actual influence under multiple edge edits. Figure˜4 presents scatter plots of predicted versus actual influence when inserting different numbers of edges. The experiments are conducted on the Cora dataset, with validation loss used as the evaluation metric. Specifically, we report results for simultaneous insertions of 10, 20, and 100 edges.
Our influence function maintains a high correlation with the actual influence, achieving 0.84 even under 100 simultaneous edge edits. Nonetheless, Figure˜4 shows that the accuracy of influence estimation diminishes as the number of simultaneous edits increases. We attribute this degradation to first-order approximation errors induced by substantial parameter shifts, a phenomenon that has also been consistently observed in prior studies on group influence estimation [22, 5, 30].
@@ 5 Applications
Table˜2 presents the adversarial attack performance of our methods, comparing it to previous influence-function-based methods and other adversarial attack methods, DICE [34] and PRBCD [14], in the white-box scenario. As a result, the GIF shows poor performance, while our method with a validation loss metric outperforms the others. An interesting observation is that, among the three metrics we tested, the validation loss-based attack is more effective than the tailored attack methods.
@@
We demonstrate the influence of edges connecting the node with the same label (homophilic edge) and different labels (heterophilic edge) on GCN. Figure˜5 presents a dumbbell plot showing the mean influence of homophilic and heterophilic edges across six datasets, which include three homophilic graphs (Cora, CiteSeer, and PubMed), where edges tend to connect nodes with the same labels, and three heterophilic graphs (Chameleon, Actor, and Squirrel), where edges tend to connect nodes with different labels.
@@
Analysis of the effect of edge rewiring
(a) BORF [26]
(b) FoSR [19]
Figure 6: Histograms of the estimated influence of edge insertions (blue) and deletions (red) selected by BORF [26] and FoSR [19], evaluated on a four-layer GCN trained on the Texas dataset. Influences are measured across validation loss, over-squashing, and over-smoothing. Arrows ($\downarrow/\uparrow$) indicate the desired direction of each measurement (decrease/increase).
Our approach provides multiple analytical perspectives on existing rewiring strategies through various evaluation metrics. Specifically, we analyze BORF [26] and FoSR [19]. BORF alleviates over-squashing by inserting edges between nodes with negative curvature and mitigates over-smoothing by removing edges with positive curvature, while FoSR alleviates over-squashing by inserting edges that enlarge the spectral gap.
Figure˜6 presents the estimated influence of edges selected by BORF and FoSR from three measurement perspectives. Edge insertions chosen by both methods generally increase the over-squashing measurement, while edge deletions in BORF tend to increase the over-smoothing measurement, confirming that each method effectively targets its intended GNN challenge. Nevertheless, we also observe unintended side effects: edge insertions often exacerbate over-smoothing, and neither insertions nor deletions consistently reduce validation loss. Mitigating these side effects could substantially enhance the overall effectiveness and reliability of edge rewiring methods.
@@ 7 Conclusion
We introduce an enhanced graph influence function that estimates the impact of edge perturbations on model predictions in Graph Neural Networks. Unlike existing graph influence functions, our approach explicitly incorporates message propagation effects and relaxes the convexity assumption, enabling it to be applied to commonly used non-convex GNNs. Additionally, we extend the framework to handle both edge deletions and insertions, broadening its applicability to real-world graph rewiring tasks. Experimental results on various real-world datasets demonstrate that our method provides significantly more accurate estimates of influence than previous methods and enables effective improvements in key target measurements such as validation loss, over-squashing, and over-smoothing. Despite these advances, our method has limitations. Although it achieves high accuracy under multiple edge edits, performance gradually declines as the number of simultaneous edits grows, and scalability to deep GNNs remains challenging. Future work could address these issues to enhance the applicability of the method to complex graph learning tasks.
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (RS-2024-00337955; RS-2023-00217286) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2024-00457882, National AI Research Lab Project; RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)).
@@ Appendix A Derivation of the influence function
In this section, we derive the influence function for linear models [21], as shown in Equation˜3, and for neural networks [3], as shown in Equation˜7. In particular, we reformulate the derivation by Bae et al. [3] to fit our setting.
@@ A.1 Influence function for linear models
We begin by considering the re-weighted objective introduced in Equation˜2:
@@
Under these conditions, the Implicit Function Theorem ensures that $\theta_{x^{\prime},y^{\prime},\epsilon}^{*}$ is continuously differentiable with respect to $\epsilon$, and we differentiate Equation˜16 using the chain rule:
@@ A.2 Influence function for neural networks
$0$ $=-\frac{1}{N}\sum_{(x,y)\in\mathcal{D}_{\text{train}}}\mathbf{J}_{h^{\theta}_{x}\theta_{s}}^{\top}\nabla_{h^{\theta}_{x}}^{2}\mathcal{L}(h^{\theta_{s}}_{x})\mathbf{J}_{h^{\theta}_{x}\theta_{s}}(\theta^{*}_{\text{lin},x^{\prime},y^{\prime},\epsilon}-\theta_{s})+\lambda(\theta^{*}_{\text{lin},x^{\prime},y^{\prime},\epsilon}-\theta_{s})+\epsilon\,\nabla_{\theta}\mathcal{L}(x^{\prime},\theta_{s}).$ (24)